# On the Robustness of Age for Learning-Based Wireless Scheduling in Unknown Environments

### v1

@@ I Introduction
Figure 1: An agricultural remote monitoring system.
@@
(iv) In Section V, we empirically investigate what happens when the i.i.d. assumption is violated, and demonstrate our key observation that the head-of-line age is more robust under periods of temporary constraint infeasibility compared to queue-length-based approaches. Another interesting result here is that the age-based policy maintains stable AoI under temporarily infeasible constraints, even compared to algorithms that attempt to optimize this metric (e.g. [25]). We also include experiments using a real data trace in this section. Finally, in Section VI, we conclude the paper.
@@ II System Model
II-A Slater’s Condition
We make the common assumption of Slater’s condition, where the throughput constraint is satisfied according to a tightness parameter $\gamma\in(0,1]$ by a $\gamma$-tight” static policy $\sigma^{\gamma}$. Let $A_{t}^{\gamma}\sim\sigma^{\gamma}$ in each timeslot $t$ such that
@@ III-B Throughput for Any Policy in Terms of Head-of-Line Age
We can also simply rewrite (7) by decomposing each set into t1t-1 disjoint subsets, i.e. $\mathcal{Q}_{k,t}={\bigcup_{\tau=1}^{t-1}\{\ell:\uptau_{k}(\ell)=\tau\}}\setminus{\bigcup_{\tau=1}^{t-1}\{\ell:d_{c}(\ell)=\tau\}}$. It follows from the definitions (9) and (10) that $|\mathcal{Q}_{k,t}|=\sum_{\tau=1}^{t-1}B_{k,\tau}-\sum_{\tau=1}^{t-1}D_{k,\tau}$, and from (15) that $|\mathcal{Q}_{k,t}|=\sum_{\tau=\uptau_{k}(\ell_{k,t}^{\to})}^{t-1}B_{k,\tau}$.
@@ III-C Throughput Violation Bounds for Age-Based Policies
$\chi_{k}-\mathop{\mathbb{E}}\left[\overline{R}_{k,t}[W]\right]\leq\frac{1+(\mathop{\mathbb{E}}\left[Z_{k,t+1}^{\to}\right]-1)(\chi_{k}+\varepsilon)}{W}-\varepsilon.$ (18)
@@ IV-A Empirical Performance
(iii) A combined queue-length-plus-TSLR-based policy similar to the algorithm found in [25], which we call “QLen+TSLR” in simulation figures, and which plays $A_{t}\in\operatorname*{argmax}_{a\in\mathcal{A}}\sum_{k=1}^{K}\left(|\mathcal{Q}_{k,t}|+\alpha\mathbb{T}_{k,t}+\eta U_{k,t}\right)I_{k}(a)$}. Note that a new tuning parameter $\alpha>0$ is introduced here.
@@
Fig. 3 shows that our age-based policy achieves middle-of-the-pack performance compared to the existing policies. Note that the TSLR-based policy achieves negative regret because it can actually achieve $\widetilde{O}(\sqrt{T}+\frac{T}{\eta})$ regret when compared against the best combinatorial arm in the unconstrained problem [11]. However, it fails to achieve the throughput requirements in general. We note that none of the policies considered seem to dominate empirically in all three performance metrics. However, in the next section, we show that the age-based policy is superior when the i.i.d. assumptions are violated and the channel conditions change abruptly.
@@ IV-B Theoretical Analysis
The theoretical results for the age-based bandit learning policy depend heavily on two Lyapunov functions: (i) the weighted quadratic Lyapunov function $L_{t}\triangleq{\frac{1}{2}\sum_{k=1}^{K}\frac{(\chi_{k}+\varepsilon)}{\overline{x}_{k}}\left(Z_{k,t}^{\to}\right)^{2}}$}, which is inspired by the Lyapunov function used in [17] and the Lyapunov function used in [25], and and (ii) the Euclidean norm of the weighted head-of-line ages $\widetilde{L}_{t}\triangleq\|(\sqrt{(\chi_{k}+\varepsilon)/\overline{x}_{k}}Z_{k,t}^{\to})_{k=1}^{K}\|_{2}$. These two Lyapunov functions are connected as follows. Since $\sqrt{x}-\sqrt{y}\leq\frac{x-y}{2\sqrt{y}}$ for all $x\geq 0$ and $y>0$, we have
@@
$\hskip-9.95863ptf(\mathbf{y})=\min_{0<\theta\leq h(\mathbf{y})}\left[\frac{9g(\mathbf{y},\theta)}{\theta^{2}}\log\left(\frac{32g(\mathbf{y},\theta)}{\theta^{2}}\right)\right]$ (23)
$\text{with }\textstyle g(\mathbf{y},\theta)=\prod_{k=1}^{K}\frac{y_{k}\,e^{2\theta/\overline{x}_{k}}}{1-(1-y_{k})e^{2\theta/\overline{x}_{k}}}$
$\text{and }\textstyle h(\mathbf{y})=-\frac{\overline{x}_{\min}}{2}\log\left(1-\min_{k\in[K]}y_{k}\right).$
@@
We would also like to point out in the following remark that the age-based bandit learning policy can achieve a data freshness guarantee of $\mathbb{T}_{k,t}=O(\eta)$ for all $k$ and $t$, although it isn’t specifically designed to optimize the TSLR $\mathbb{T}_{k,t}$. It’s also worth noting that queue-length-based policies (e.g. [12]) can also achieve similar guarantees (see Lemma 1 in [25]), but recall that the age-based policy achieves empirically better performance on this metric (see Section IV-A).
@@
$\mathrm{Reg}_{T}$ $\leq\frac{\varepsilon KT}{\gamma}+\frac{T}{\eta}\sum_{k=1}^{K}\frac{1}{\chi_{k}\overline{x}_{k}}+\underbrace{\sqrt{56KI_{\max}T\log T}+K\pi^{2}/3}_{\text{from standard UCB regret analysis}}$ (26)
@@
$\hskip-9.95863pt\phantom{\,=\,}+\sum_{t=1}^{T}\sum_{k=1}^{K}\mathop{\mathbb{E}}\left[I_{k}(A_{t})\,U_{k,t}-R_{k}(S_{t},A_{t})\right]+\frac{T}{\eta}\sum_{k=1}^{K}\frac{1}{\chi_{k}\overline{x}_{k}}.$
@@ V-A Experimental Setup
(a) Short-Term Throughput $\overline{R}_{k,t}(W)$
@@
Figure 5: Operation of the queue-length-based policy vs. our age-based policy (19) when the throughput constraint for arm $k=1$ is infeasible. Here, one simulation run from the results in Fig. 4 is shown.
@@ V-B Discussion of Results
Fig. 4 shows the result of these experiments and we use the same algorithm parameters from Section IV-A for consistency. In this figure, dotted lines are associated with arm $k=1$ and solid lines are associated with arm $k=2$. The key observation is that our age-based policy (shown in blue) is able to quickly recover from the period of temporary constraint infeasibility (the shaded region). Focusing first on the short-term throughput in Fig. 4(a), the two queue-length-based policies (“QLen” and “QLen+TSLR”) neglect arm 2 and continue trying to schedule arm 1, whose queue length is blowing up. On the other hand, our age-based policy (19) never completely neglects arm 2. After the period of constraint infeasibility, the queue-length-based policies never recover. Interestingly, our age-based policy, which is not explicitly designed to target the TSLR metric, does much better than the QLen+TSLR policy on this metric in the presence of constraint infeasibility. Our intuition is that for QLen+TSLR, the queue length metric completely dominates any other metric considered in the algorithm. This also demonstrates the utility of the $O(\eta)$ TSLR control afforded by our age-based policy (see Remark 2). As expected, the TSLR-based policy does well in the TSLR metric, but completely blunders the short-term throughput since it is not designed for that. For the TSLR-based policy, the opposite changes in short-term throughput during the period of constraint infeasibility are due to the changes in reward between the arms.
To shed some light on why the head-of-line age does not suffer as much from periods of constraint infeasibility compared to the queue length, we plot the operation of the queue-length-based policy and our age-based policy for one simulation run in Fig. 5. Since we consider one action per arm, the “weight” on an action $a=k$ is $\eta U_{k,t}+Z_{k,t}^{\to}$ under our age-based policy (19) and $\eta U_{k,t}+|\mathcal{Q}_{k,t}|$ under the queue-length-based policy. The key difference is in the dynamic of the head-of-line age $\Delta Z_{k,t}^{\to}$ given by Lemma 2 and the dynamic of the queue length $\Delta|\mathcal{Q}_{k,t}|=B_{k,t}-D_{k,t}$. In particular, the head-of-line age always increments by 1 when a reward is not obtained, and therefore the head-of-line age for the neglected arm always has a chance to “catch up” to the other arm. On the other hand, the queue length grows much more slowly at a rate of $\chi_{k}+\varepsilon$. From these results, we argue that this robustness under constraint infeasibility incentivizes the use of the head-of-line age in place of the virtual queue length where appropriate in drift-plus-penalty style algorithms deployed to real-world systems.
@@ V-C Replication of the Experiment Using a Real Data Trace
(a) Short-Term Throughput $\overline{R}_{k,t}(W)$

### v2

@@ I Introduction
Figure 1: An agricultural remote monitoring system with short-term throughput requirements and abruptly changing wireless channel conditions.
@@
(iv) In Section V, we empirically investigate what happens when the i.i.d. assumption is violated, and demonstrate our key observation that the head-of-line age is more robust under periods of temporary constraint infeasibility compared to queue-length-based approaches. Another interesting result here is that the age-based policy maintains stable TSLR under temporarily infeasible constraints, even compared to algorithms that attempt to optimize this metric (e.g. [26]). We also include experiments using a real data trace in this section. Finally, in Section VI, we conclude the paper.
@@
Proofs: All omitted proofs of theoretical results can be found in the Appendix.
@@ II System Model
Finally, we make the common assumption of Slater’s condition [^3] This assumption may potentially be removed using techniques from [6]., where the throughput constraint is satisfied according to a tightness parameter $\gamma\in(0,1]$ by a $\gamma$-tight” static policy $\sigma^{\gamma}$. Let $A_{t}^{\gamma}\sim\sigma^{\gamma}$ denote the action under the $\gamma$-tight static policy in each timeslot $t$, which is assumed to satisfy
@@ III-B Throughput for Any Policy in Terms of Head-of-Line Age
We can also simply rewrite (7) by decomposing each set into t1t-1 disjoint subsets, i.e. $\mathcal{Q}_{k,t}={\bigcup_{\tau=1}^{t-1}\{\ell:\uptau_{k}(\ell)=\tau\}}\setminus{\bigcup_{\tau=1}^{t-1}\{\ell:d_{k}(\ell)=\tau\}}$. It follows from the definitions (9) and (10) that $|\mathcal{Q}_{k,t}|=\sum_{\tau=1}^{t-1}B_{k,\tau}-\sum_{\tau=1}^{t-1}D_{k,\tau}$, and from (15) that $|\mathcal{Q}_{k,t}|=\sum_{\tau=\uptau_{k}(\ell_{k,t}^{\to})}^{t-1}B_{k,\tau}$.
@@ III-C Throughput Violation Bounds for Age-Based Policies
$\chi_{k}-\mathop{\mathbb{E}}\left[\overline{R}_{k,t}[W]\right]\leq\frac{1+\left(\mathop{\mathbb{E}}\left[Z_{k,t+1}^{\to}\right]-1\right)(\chi_{k}+\varepsilon)}{W}-\varepsilon.$ (18)
@@ IV-A Empirical Performance
(iii) A combined queue-length-plus-TSLR-based policy similar to the algorithm found in [26], which we call “QLen+TSLR” in simulation figures, and which plays $A_{t}\in\operatorname*{argmax}_{a\in\mathcal{A}}\sum_{k=1}^{K}\left(|\mathcal{Q}_{k,t}|+\alpha\mathbb{T}_{k,t}+\eta U_{k,t}\right)I_{k}(a)$}. Note that a new tuning parameter $\alpha$ is introduced here for the TSLR term.
@@
Fig. 3 shows that our age-based policy achieves similar performance to the queue-length-based policy in this i.i.d. environment. Note that the TSLR-based policy achieves negative regret because it can actually achieve $\widetilde{O}(\sqrt{T}+\frac{T}{\eta})$ regret when compared against the best combinatorial arm in the unconstrained problem [12]. However, it fails to achieve the throughput requirements in general. We note that none of the policies considered seem to dominate empirically in all three performance metrics. However, in the next section, we show that the age-based policy is superior in the non-i.i.d. case where the channel conditions change abruptly.
@@ IV-B Theoretical Analysis
The theoretical results for the age-based bandit learning policy depend heavily on two Lyapunov functions: (i) the weighted quadratic Lyapunov function $L_{t}\triangleq{\frac{1}{2}\sum_{k=1}^{K}\frac{(\chi_{k}+\varepsilon)}{\overline{x}_{k}}\left(Z_{k,t}^{\to}\right)^{2}}$}, which is inspired by the Lyapunov function used in [18] and the Lyapunov function used in [26], and (ii) the Euclidean norm of the weighted head-of-line ages $\widetilde{L}_{t}\triangleq\|(\sqrt{(\chi_{k}+\varepsilon)/\overline{x}_{k}}Z_{k,t}^{\to})_{k=1}^{K}\|_{2}$. These two Lyapunov functions are connected as follows. Since $\sqrt{x}-\sqrt{y}\leq\frac{x-y}{2\sqrt{y}}$ for all $x\geq 0$ and $y>0$, we have
@@
$f(\mathbf{y})=\min_{0<\theta\leq h(\mathbf{y})}\left[\frac{9g(\mathbf{y},\theta)}{\theta^{2}}\log\left(\frac{32g(\mathbf{y},\theta)}{\theta^{2}}\right)\right]$ (23)
where $h(\mathbf{y})=-\frac{\overline{x}_{\min}}{2}\log\left(1-\min_{k\in[K]}y_{k}\right)$, $g(\mathbf{y},\theta)=\prod_{k=1}^{K}\frac{y_{k}\,e^{2\theta/\overline{x}_{k}}}{1-(1-y_{k})e^{2\theta/\overline{x}_{k}}}$, and $\overline{x}_{\min}\triangleq\min_{k}\overline{x}_{k}$.
@@
We would also like to point out in the following remark that the age-based bandit learning policy can achieve a data freshness guarantee of $\mathbb{T}_{k,t}=O(\eta)$ for all $k$ and $t$, although it isn’t specifically designed to optimize the TSLR $\mathbb{T}_{k,t}$. It’s also worth noting that queue-length-based policies (e.g. [13]) can also achieve similar guarantees (see Lemma 1 in [26]).
@@
$\mathrm{Reg}_{T}$ $\leq\frac{\varepsilon KT}{\gamma}+\frac{T}{\eta}\sum_{k=1}^{K}\frac{1}{\chi_{k}\overline{x}_{k}}+\underbrace{\sqrt{56KI_{\max}T\log T}+\frac{K\pi^{2}}{3}}_{\text{from standard UCB regret analysis}}$ (26)
@@
$\hskip-5.69046pt+\sum_{t=1}^{T}\sum_{k=1}^{K}\mathop{\mathbb{E}}\left[I_{k}(A_{t})\,U_{k,t}-R_{k}(S_{t},A_{t})\right]+\frac{T}{\eta}\sum_{k=1}^{K}\frac{1}{\chi_{k}\overline{x}_{k}}.$
@@ V-A Experimental Setup
(a) Short-Term Throughput $\overline{R}_{k,t}[W]$
@@
Figure 5: Operation of the queue-length-based policy vs. our age-based policy (19) when the throughput constraint for arm $k=1$ is infeasible. Here, one simulation run from the results in Fig. 4 is shown. Vertical dotted red lines show when action 2 is played, and action 1 is played in all other timeslots.
@@ V-B Discussion of Results
Fig. 4 shows the result of these experiments and we use the same algorithm parameters from Section IV-A for consistency. In this figure, dotted lines are associated with arm $k=1$ and solid lines are associated with arm $k=2$. The key observation is that our age-based policy (shown in blue) is able to quickly recover from the period of temporary constraint infeasibility (the shaded region). Focusing first on the short-term throughput in Fig. 4(a), the two queue-length-based policies (“QLen” and “QLen+TSLR”) neglect arm 2 and continue trying to schedule arm 1, whose queue length is blowing up. On the other hand, our age-based policy (19) never completely neglects arm 2. After the period of constraint infeasibility, the queue-length-based policies never recover. Interestingly, our age-based policy, which is not explicitly designed to target the TSLR metric, does much better than the QLen+TSLR policy on this metric, which never recovers from the period of constraint infeasibility. Our intuition is that for QLen+TSLR, the queue length metric completely dominates any other metric considered in the algorithm. This also demonstrates the utility of the $O(\eta)$ TSLR control afforded by our age-based policy (see Remark 2). As expected, the TSLR-based policy does well in the TSLR metric, but completely blunders the short-term throughput since it is not designed for that. For the TSLR-based policy, the opposite changes in short-term throughput during the period of constraint infeasibility are due to the changes in reward between the arms.
To shed some light on why the head-of-line age does not suffer as much from periods of constraint infeasibility compared to the queue length, we plot the operation of the queue-length-based policy and our age-based policy for one simulation run in Fig. 5. Since we consider one action per arm, the “weight” on an action $a=k$ is $\eta U_{k,t}+Z_{k,t}^{\to}$ under our age-based policy (19) and $\eta U_{k,t}+|\mathcal{Q}_{k,t}|$ under the queue-length-based policy. The key difference is in the dynamic of the head-of-line age $\Delta Z_{k,t}^{\to}$ given by Lemma 2 and the dynamic of the queue length $\Delta|\mathcal{Q}_{k,t}|=B_{k,t}-D_{k,t}$. In particular, the head-of-line age always increments by 1 when a reward is not obtained, and therefore the head-of-line age for the neglected arm always has a chance to “catch up” to the other arm. On the other hand, the queue length grows much more slowly at a rate of $\chi_{k}+\varepsilon$. From these results and the fact that the age-based and queue-length-based policies perform similarly in the i.i.d. setting (see Section IV-A), we argue that this robustness under constraint infeasibility incentivizes the use of the head-of-line age in place of the queue length where appropriate in drift-plus-penalty style algorithms deployed to real-world systems.
@@ V-C Replication of the Experiment Using a Real Data Trace
(a) Short-Term Throughput $\overline{R}_{k,t}[W]$